Solve, write your answer in interval notation and graph the solution set.

14c. |2x-1| + 7 greater than or equal to 1

14d. |5x+3| less than or equal to -1
1
Expert's answer
2013-01-23T11:28:01-0500
14c. |2x-1| + 7 greater than or equal to 1 |2x-1| + 7>=1
|2x-1|>=-6
but |t|>=0>-6 for every t
so solution Real line (the whole set of real numbers)

14d. |5x+3| lessthan or equal to -1
|5x+3|>=0 for every x so solution - empty set